# Diversity of rocky planet atmospheres in the C-H-O-N-S-Cl system with interior dissolution, non-ideality, and condensation: Application to TRAPPIST-1e and sub-Neptunes

### v1

@@ 3.1.2 Volatile partitioning during the magma ocean stage
We consider TRAPPIST-1e with a fully molten mantle, for which solubility and gas phase equilibria are both important in determining the atmospheric mass and composition. For comparison, Appendix A presents the same analysis assuming a partially molten mantle with 10% melt fraction. For a fully molten mantle, solubility conspires to restrict the diversity of atmospheres (Figure 1) compared to the low melt fraction case (Figure A1). The data are coloured according to the oxygen fugacity relative to the IW buffer (IW), which defines whether the atmosphere is reduced (lower values) or oxidised (higher values). The scatter plots capture the diversity of atmospheres, whereas the contoured density levels and marginal distributions reveal the most probable range of values. Figure 1 and Figure 2 show major and minor species, respectively.
@@
A fully molten mantle always dissolves at least 60% of the total hydrogen budget. At IW close to , high C/H atmospheres result in marginally less total H dissolution, all else being equal, due to the stability of CH4(g) over H2(g). For all but these most reduced conditions, the dissolved fraction of H approaches 100% (Figure 4). This is evident by the sharp peak around 100% in the marginal distribution of dissolved hydrogen where even the 10th percentile exceeds 97% (see also Sossi et al., 2020). As a result, atmospheric C/H ratios can be around 4 orders of magnitude larger than total C/H, with the disparity increasing to more oxidised conditions due to increasing H2O/H2. Nearly no C is dissolved in the mantle under reducing conditions (owing to the low solubilities of CO(g) and CH4(g)), but this figure increases up to about 25% of the total carbon inventory at the most oxidised conditions where CO2(g) is stable (Figure 4). This proportion decreases for correspondingly lower melt fractions (Figure A3).
@@
Total pressures range between 100 to 104 bar (90% lie within 5–500 bar), with atmospheres produced in equilibrium at intermediate O2s near the IW buffer delimiting both the high- and low ends of the range. The total pressure is largely governed by the number of ocean masses of H, together with the C/H ratio, with high C/H resulting in the highest atmospheric pressures at a given O2. Even though the mantles in equilibrium with reducing atmospheres store proportionally less H and C than their more oxidising counterparts, their lower mean molecular masses () result in only modest total pressures (Equation E1). At high O2s near IW=+4, the increasing degree of CO2 dissolution into the mantle restricts the total pressure of SO2-rich atmospheres to around 200 bar. Sulfur is almost entirely dissolved below IW=+3, but above this value up to 60% of the sulfur inventory can exist in the atmosphere (Figure 4). The non-monotonic solubility of S arises due to the occurrence of two moieties in silicate liquids; S2- below IW=+4 and SO above this value (O’Neill & Mavrogenes, 2022; Boulliung & Wood, 2023). Because the S content of the silicate liquid is proportional to O (for constant temperature, S2 and FeO content) when S2- is stable, and to O when SO is stable, the solubility of S exhibits a minimum at S2-/SO = 0.5 (where is mole fraction), or roughly IW=+4 (see also Hughes et al., 2023). Therefore, SO2 should decrease in atmospheres that are more oxidising and more reducing than those formed at the solubility minimum.
@@
Minor species for a 100% molten mantle are shown in Figure 2 (see Figure A2 for the 10% molten equivalent). Of these, N2(g) can become a major species (i.e., its mixing ratio exceeds 0.5) in a handful of cases. These atmospheres correspond to those with the lowest total pressures. That is, since the abundances of N, Cl and S are held constant (while those of H, C and O vary) in the Monte Carlo simulations, N2-rich atmospheres arise when the prescribed initial abundances of H and C are low. This behaviour is also reflected in the inverse correlation between Cl2(g) mixing ratio (which never exceeds 1.510-7) and total pressure (Figure 2). Indeed, the mixing ratios of N2 and Cl2 are positively correlated. The other minor species diverge from this expectation, because their partial pressures are sensitive to the presence of other volatile species, whose fugacities depend on O2 and/or the bulk chemical composition of the atmosphere.
@@
The second most abundant minor species is S2(g), reaching mixing ratios up to 3% (0.03), with these cases corresponding exclusively to oxidising atmospheres. As detailed above, more reducing conditions promote the dissolution of S as S2- in the silicate liquid, thereby restricting the S budget of the atmosphere and thus its ability to stabilise S2(g). A third S-bearing gas species, H2S(g), is less abundant than S2(g) and SO2(g) in atmospheres more oxidised than IW=+2, yet it becomes the dominant S-bearing gas species (mixing ratios up to 0.005) at O2s below this value. This reflects the increase in H2 at the expense of H2O, promoting the formation of H2S by:
@@ 3.1.3 Cooled atmosphere and elemental surface inventory
The magma ocean stage establishes the initial volatile inventories in the atmosphere (Section 3.1.2), and as the atmosphere cools and condenses it distributes materials at the planetary surface that can participate in geochemical cycles and fuel a potential biosphere. To this end, we begin with the calculated atmospheric elemental abundances for each of the fully molten mantle cases presented in Section 3.1.2 and cool the atmospheres, isochemically (i.e., assuming a constant bulk atmosphere composition inherited from the magma ocean stage), to 280 K. This is a few kelvin above the freezing point of water at 1 bar and lies between the estimates of the equilibrium temperatures of TRAPPIST-1d and TRAPPIST-1e assuming a null albedo (Delrez et al., 2018). We include the contribution of any graphite present in the abundance of carbon of the high temperature atmosphere, given that graphite remains close to the planetary surface (Keppler & Golabek, 2019) and in equilibrium with the atmosphere. In addition to the aforementioned gas species and graphite (Section 3.1.1), we further allow for the presence of liquid water (H2O), rhombic sulfur (also known as -sulfur), and ammonium chloride (ClH4N) depending on thermodynamic stability, since these condensed phases can be major carriers of elements in the system. The same analysis is applied to the 10% molten mantle cases in Appendix A.
These atmospheres can also produce up to 0.3 Earth-ocean equivalents of -sulfur (Figure 5). Crucially, the starting conditions that favour large amounts of sulfur precipitation (oxidised) are different from the conditions that favour significant water, graphite, and ammonium chloride precipitation (IW and below). This arises because S is more abundant in the high-temperature atmosphere at oxidising conditions, while C and particularly H are more abundant at more reducing conditions (see Section 3.1.2, Figure 4). There is also a trade-off between graphite and water, since high C/H in the starting conditions promotes larger graphite inventories whereas low C/H forms larger water oceans. In the family of atmospheres with the lowest molar masses () produced by initially low C/H ratios and high H mass fractions, graphite saturation is not reached. The median total atmospheric pressure is roughly 65 bar and the maximum is a few kbar for oxidised starting conditions with high C/H and high initial number of water oceans.
Consideration of cases in which the mantle is only 10% molten leads to cooled atmospheres that have greater mixing ratios of H-bearing species. This leads to a more continuous distribution in across the simulation space (Figure A4), in which the heights of peaks associated with H-bearing species (notably H2(g) and CH4(g), given that H2O is always condensed) compared to that of CO2(g) are relatively greater than for the 100% molten cases. Nevertheless, the four families of atmospheres previously identified for the fully molten case are again evident, albeit slightly less distinct. CO2 atmospheres are similarly dominant, with a tail extending to SO2 atmospheres. Table 3 also reveals the formation of H2S-rich atmospheres for the molten case, although they are low in number (2% of the total).
3.2 K2-18b and similar sub-Neptune-sized planets
3.2.1 Context and parameters
Many sub-Neptunes likely harbour massive magma oceans beneath large insulating atmospheres, a configuration that corresponds to so-called ‘gas dwarfs’. Such a structure has been proposed for the canonical sub-Neptune K2-18b (Shorttle et al., 2024; Wogan et al., 2024), although an alternative hypothesis suggests it might support clement conditions at the base of its atmosphere (Madhusudhan et al., 2023). Other sub-Neptune-sized exoplanets are currently being observed with JWST (e.g., TOI-270 d, GJ 1214 b), and preliminary observations reveal absorption features of various gas species (e.g., H2O, CO2, CH4) (Kempton et al., 2023; Benneke et al., 2024). However, whether or not these planets possess a magma ocean surface, a temperate surface that could support liquid water, or no defined surface at all, remains unknown. Distinguishing between these scenarios requires self-consistent modelling frameworks that adequately capture the governing physics and chemistry of both the interior and the atmosphere. To this end, Atmodeller allows testing the gas dwarf model with an important addition that has often been neglected in previous work on sub-Neptunes (Schlichting & Young, 2022; Misener et al., 2023; Young et al., 2023; Shorttle et al., 2024; Rogers et al., 2024; Rigby et al., 2024), though see Tian & Heng (2024); Seo et al. (2024): the role of non-ideality, in which the volatile species at the conditions of the magma ocean-atmosphere interface do not obey the ideal gas law. Rather, real gases decouple their partial pressure (which is related to molar abundance, mass balance, and mechanical pressure) from their thermodynamic pressure which regulates equilibrium chemistry and solubility (Section 2.3).
We simulate K2-18 b as a gas dwarf and demonstrate the effects of non-ideality on sub-Neptune atmospheres. K2-18b has a total mass of kg (8.63 M⊕) and an estimated mantle mass of kg (assuming the same core/mantle mass ratio as the Earth) (Benneke et al., 2019). K2-18b’s surface radius, assuming it has an optically thick atmosphere overlying a magma ocean surface, is estimated to be 11225 km (1.76 R⊕; assuming an Earth-like interior composition and the super-Earth equation of state taken from Table 2 of Hakim et al., 2018). The surface temperature at the magma ocean–atmosphere interface is set at 3000 K, which is at or above the upper limit of the calibration for current solubility laws and real gas EOS, and the mantle melt fraction is set to 100%. We consider 5 gas species in the atmosphere, each with their corresponding solubility laws, necessarily assuming the mantle composition is basaltic to remain internally consistent: H2O (Dixon et al., 1995), H2 (Hirschmann et al., 2012), CO (Yoshioka et al., 2019), CO2 (Dixon et al., 1995), and CH4 (Ardia et al., 2013). O2 is also included as a gas species. For the cases with real gases, the Compensated Redlich-Kwong (CORK) model is used for H2O, an ab-initio EOS model for H2 (Chabrier & Debras, 2021), and a corresponding states model for CO, CO2, and CH4 (Shi & Saxena, 1992) (Table 2). O2 is assumed to behave as an ideal gas in all cases, which does not introduce any uncertainty owing to its vanishingly low fugacities under the range of conditions studied.
To explore the effects of non-ideality, we perform two series of simulations in which volatiles can partition between the silicate mantle and the atmosphere. The total hydrogen inventory is varied from 0.1 to 3% of K2-18b’s mass (Figures 6 and 7) assuming both real and ideal gas behaviour. In these simulations, all other parameters are held constant: K, C/H (by mass) = 0.32 ( the solar ratio, as recently inferred for K2-18b, Wogan et al. (2024)) and O2 at IW=-3. The equivalent cases, in which each of the other two variables (IW and C/H ratio) are varied while the others are held constant, are shown in Figures F1 and F2, respectively. All simulation data are available to download (Bower, 2025).
3.2.2 Volatile storage in the mantle
For low H mass fractions, CO(g) is the dominant species, but is exceeded by H2(g) across all hydrogen mass fractions 0.2% in the ideal case, whereas CH4 exceeds H2 above 0.3% H mass fraction in the real case. The predominance of CH4(g) results from the interplay between total pressure and the fugacity coefficients of the major atmospheric species, as described by the homogeneous equilibrium:
where, at equilibrium, and at constant O2:
where is the total mass of the atmosphere, is the acceleration due to gravity, and is the surface radius. Because pressure is proportional to 1/, even small changes in surface radius have tangible effects on total pressure, for a constant atmospheric mass budget. Conversely, the linear dependence of on (and hence on the planetary mass, ) results in a more modest mutual dependence. However, because solubility is included, is not constant, and increasing pressure for a constant total mass budget results in higher dissolution (since , where is any given gas species), giving rise to the monotonic decrease in H2 (and hence total ) with increasing or with decreasing in the ideal case. Conversely, increasing leads to increases in (analogous to the mechanism discussed with respect to Equation 6, above), promoting additional H dissolution as molecular H2. The net result is sub-Neptunes are expected to be CH4-rich for cases in which mean densities are roughly 6500 kg/m3 at IW=-3, H=3% , and bulk C/H=100 solar. Figures F1 and F2 illustrate the effects of varying oxygen fugacity from (at constant C/H = 0.32 and H mass fraction = 1% ) and total C/H mass ratio (at constant O2 = IW=-3 and H mass fraction = 1% ), respectively, on volatile speciation in the atmosphere and dissolution in the interior assuming real and ideal gas behaviour, for K2-18b-like cases.
4 Discussion
4.1 Earth-sized planets
For Earth-sized planets, our results show that the dominant atmosphere at temperate conditions that emerges after planet formation is CO2 rich with a total surface pressure around 100 bar. Such an atmosphere forming upon cooling on the early Earth from a magma ocean state was first hypothesised by Holland (1984), computed by Zahnle et al. (2007) and demonstrated by Sossi et al. (2020) for a magma ocean initially equilibrated at its estimated O2 close to the IW buffer. CO2-rich atmospheres are compatible with that of Venus at the present day, both in composition and total pressure. The composition is also similar to that of modern Mars, hinting that this mode of atmosphere formation may have been commonplace on rocky planets in our solar system. That the Earth does not currently possess a 97:3 CO2–N2 atmosphere points to the operation of geochemical cycles that drew down carbon from the atmosphere through weathering by the Urey reaction to form carbonates (Sleep & Zahnle, 2001; Zahnle et al., 2007) that are subsequently recycled into the interior, thus evolving its atmosphere away from CO2-rich over geological timescales. This is because carbon is abundant, CO2 is strongly favoured at 280 K, and water condensation buffers the emergence of hydrogen-dominated atmospheres even when total C/H is low (as with atmospheres derived from a partially molten starting state, Appendix A).

### v2

@@ 3.1.2 Volatile partitioning during the magma ocean stage
We consider TRAPPIST-1e with a fully molten mantle, for which solubility and gas phase equilibria are both important in determining the atmospheric mass and composition. For comparison, Appendix B presents the same analysis assuming a partially molten mantle with 10% melt fraction. For a fully molten mantle, solubility conspires to restrict the diversity of atmospheres (Figure 1) compared to the low melt fraction case (Figure B1). The data are coloured according to the oxygen fugacity relative to the IW buffer (IW), which defines whether the atmosphere is reduced (lower values) or oxidised (higher values). The scatter plots capture the diversity of atmospheres, whereas the contoured density levels and marginal distributions reveal the most probable range of values . Figure 1 and Figure 2 show major and minor species, respectively.
@@
A fully molten mantle always dissolves at least 60% of the total hydrogen budget. At IW close to high C/H atmospheres have marginally less total H dissolved, all else being equal, due to the stability of CH4(g) over H2(g). For all but these most reduced conditions, the dissolved fraction of H approaches 100% (Figure 4). This is evident by the sharp peak around 100% in the marginal distribution of dissolved hydrogen where even the 10th percentile exceeds 97% (see also Sossi et al., 2020). As a result, atmospheric C/H ratios can be around 4 orders of magnitude larger than total C/H, with the disparity increasing to more oxidised conditions due to increasing H2O/H2. Nearly no C is dissolved in the mantle under reducing conditions (owing to the low solubilities of CO(g) and CH4(g)), but this value increases up to about 25% of the total carbon inventory at the most oxidised conditions where CO2(g) is stable (Figure 4). This proportion decreases for correspondingly lower melt fractions (Figure B3).
@@
Total pressures range between 100 to 104 bar (90% lie within 5–500 bar), with atmospheres produced in equilibrium at intermediate O2s near the IW buffer delimiting both the high- and low ends of the range. The total pressure is largely governed by the number of ocean masses of H, together with the C/H ratio, with high C/H resulting in the highest atmospheric pressures at a given O2. Even though the mantles in equilibrium with reducing atmospheres store proportionally less H and C than their more oxidising counterparts, their lower mean molecular masses () result in only modest total pressures (Equation F1). At high O2s near IW=+4, the increasing degree of CO2 dissolution into the mantle restricts the total pressure of SO2-rich atmospheres to around 200 bar. Sulfur is almost entirely dissolved below IW=+3, but above this value up to 60% of the sulfur inventory can exist in the atmosphere (Figure 4). The non-monotonic solubility of S arises due to the occurrence of two moieties in silicate liquids; S2- below IW=+4 and SO above this value (O’Neill & Mavrogenes, 2022; Boulliung & Wood, 2023). Because the S content of the silicate liquid is proportional to O (for constant temperature, S2 and FeO content) when S2- is stable, and to O when SO is stable, the solubility of S exhibits a minimum at S2-/SO = 0.5 (where is mole fraction), or roughly IW=+4 (see also Hughes et al., 2023). Therefore, SO2 should decrease in atmospheres that are more oxidising and more reducing than those formed at the solubility minimum.
@@
Minor species for a 100% molten mantle are shown in Figure 2 (see Figure B2 for the 10% molten equivalent). Of these, N2(g) can become a major species (i.e., its mixing ratio exceeds 0.5) in a handful of cases. These atmospheres correspond to those with the lowest total pressures. That is, since the abundances of N, Cl and S are held constant (while those of H, C and O vary) in the Monte Carlo simulations, N2-rich atmospheres arise when the prescribed initial abundances of H and C are low. This behaviour is also reflected in the inverse correlation between Cl2(g) mixing ratio (which never exceeds 1.510-7) and total pressure (Figure 2). Indeed, the mixing ratios of N2(g) and Cl2(g) are positively correlated. The other minor species diverge from this expectation, because their partial pressures are sensitive to the presence of other volatile species, whose fugacities depend on O2 and/or the bulk chemical composition of the atmosphere.
@@
The second most abundant minor species is S2(g), reaching mixing ratios up to 3% (0.03), with these cases corresponding exclusively to oxidising atmospheres. As detailed above, more reducing conditions promote the dissolution of S as S2- in the silicate liquid, thereby restricting the S budget of the atmosphere and thus its ability to stabilise S2(g). A third S-bearing gas species, H2S(g), is less abundant than S2(g) and SO2(g) in atmospheres more oxidised than IW=+2, yet it becomes the dominant S-bearing gas species (mixing ratios up to 0.005) at O2s below this value. This reflects the increase in H2 at the expense of H2O, promoting the formation of H2S(g) by:
@@ 3.1.3 Cooled atmosphere and elemental surface inventory
The magma ocean stage establishes the initial volatile inventories in the atmosphere (Section 3.1.2), and as the atmosphere cools and condenses it distributes materials at the planetary surface that can participate in geochemical cycles and fuel a potential biosphere. To this end, we begin with the calculated atmospheric elemental abundances for each of the fully molten mantle cases presented in Section 3.1.2 and cool the atmospheres, isochemically (i.e., assuming a constant bulk atmosphere composition inherited from the magma ocean stage), to 280 K. This is a few kelvin above the freezing point of water at 1 bar and lies between the estimates of the equilibrium temperatures of TRAPPIST-1d and TRAPPIST-1e assuming a null albedo (Delrez et al., 2018). We include the contribution of any graphite present in the abundance of carbon of the high temperature atmosphere, given that graphite remains close to the planetary surface (Keppler & Golabek, 2019) and in equilibrium with the atmosphere. In addition to the aforementioned gas species and graphite (Section 3.1.1), we further allow for the presence of liquid water (H2O), rhombic sulfur (also known as -sulfur), and ammonium chloride (ClH4N) depending on thermodynamic stability, since these condensed phases can be major carriers of elements in the system. The same analysis is applied to the 10% molten mantle cases in Appendix B.
@@
These atmospheres can also produce up to 0.3 Earth-ocean equivalents of -sulfur (Figure 5). Crucially, the starting conditions that favour large amounts of sulfur precipitation (oxidised) are different from the conditions that favour significant water, graphite, and ammonium chloride precipitation (IW and below). This arises because S is more abundant in high-temperature atmospheres at oxidising conditions, while C and particularly H are more abundant at more reducing conditions (see Section 3.1.2, Figure 4). There is also a trade-off between graphite and water, since high C/H in the starting conditions promotes larger graphite inventories whereas low C/H forms larger water oceans. In the family of atmospheres with the lowest molar masses () produced by initially low C/H ratios and high H mass fractions, graphite saturation is not reached. The median total atmospheric pressure is roughly 65 bar and the maximum is a few kbar for oxidised starting conditions with high C/H and high initial number of water oceans.
Consideration of cases in which the mantle is only 10% molten leads to cooled atmospheres that have greater mixing ratios of H-bearing species. This leads to a more continuous distribution in across the simulation space (Figure B4), in which the heights of peaks associated with H-bearing species (notably H2(g) and CH4(g), given that H2O is always condensed) compared to that of CO2(g) are relatively greater than for the 100% molten cases. Nevertheless, the four families of atmospheres previously identified for the fully molten case are again evident, albeit slightly less distinct. CO2 atmospheres are similarly dominant, with a tail extending to SO2 atmospheres. Table 3 also reveals the formation of H2S-rich atmospheres for the molten case, although they are low in number (2% of the total).
3.2 K2-18b and similar sub-Neptune-sized planets
3.2.1 Context and parameters
Many sub-Neptunes likely harbour massive magma oceans beneath large insulating atmospheres, a configuration that corresponds to so-called ‘gas dwarfs’. Such a structure has been proposed for the canonical sub-Neptune K2-18b (Shorttle et al., 2024; Wogan et al., 2024), although an alternative hypothesis suggests it might support clement conditions at the base of its atmosphere (Madhusudhan et al., 2023). Other sub-Neptune-sized exoplanets are currently being observed with JWST (e.g., TOI-270 d, GJ 1214 b), and preliminary observations reveal absorption features of various gas species . However, whether or not these planets possess a magma ocean surface, a temperate surface that could support liquid water, or no defined surface at all (e.g., Young et al., 2024b), remains unknown. Distinguishing between these scenarios requires self-consistent modelling frameworks that adequately capture the governing physics and chemistry of both the interior and the atmosphere. To this end, Atmodeller allows testing the gas dwarf model with an important addition that has often been neglected in previous work on sub-Neptunes (Schlichting & Young, 2022; Misener et al., 2023; Young et al., 2023; Shorttle et al., 2024; Rogers et al., 2024; Rigby et al., 2024), though see Tian & Heng (2024); Seo et al. (2024): the role of non-ideality, in which the volatile species at the conditions of the magma ocean-atmosphere interface do not obey the ideal gas law. Rather, real gases decouple their partial pressure (which is related to molar abundance, mass balance, and mechanical pressure) from their thermodynamic pressure which regulates equilibrium chemistry and solubility (Section 2.3).
We simulate K2-18 b as a gas dwarf and demonstrate the effects of non-ideality on sub-Neptune atmospheres. K2-18b has a total mass of kg (8.63 M⊕) and an estimated mantle mass of kg (assuming the same core/mantle mass ratio as the Earth) (Benneke et al., 2019). K2-18b’s surface radius, assuming it has an optically thick atmosphere overlying a magma ocean surface, is estimated to be 11225 km (1.76 R⊕; assuming an Earth-like interior composition and the super-Earth mass–radius relation taken from Table 2 of Hakim et al., 2018). The surface temperature at the magma ocean–atmosphere interface is set at 3000 K, We further note that this temperature is at or above the upper limit of the calibration for current solubility laws and real gas EOS. The mantle melt fraction is set to 100%. We consider 6 gas species in the atmosphere, each with their corresponding solubility laws, necessarily assuming the mantle composition is basaltic to remain internally consistent: H2O (Dixon et al., 1995), H2 (Hirschmann et al., 2012), CO (Yoshioka et al., 2019), CO2 (Dixon et al., 1995), and CH4 (Ardia et al., 2013). O2 is also included as a gas species. For the cases with real gases, the Compensated Redlich-Kwong (CORK) model is used for H2O, an ab-initio EOS model for H2 (Chabrier & Debras, 2021), and a corresponding states model for CO, CO2, and CH4 (Shi & Saxena, 1992) (Table 2). O2 is assumed to behave as an ideal gas in all cases, which does not introduce any uncertainty owing to its vanishingly low fugacities under the range of conditions studied.
To explore the effects of non-ideality, we perform two series of simulations in which volatiles can partition between the silicate mantle and the atmosphere. The total hydrogen inventory is varied from 0.1 to 3% of K2-18b’s mass (Figures 6 and 7) assuming both real and ideal gas behaviour. In these simulations, all other parameters are held constant: K, C/H (by mass) = 0.32 ( the solar ratio, as recently inferred for K2-18b, Wogan et al. (2024)) and O2 at IW=-3. The equivalent cases, in which each of the other two variables (IW and C/H ratio) are varied while the others are held constant, are shown in Figures G1 and G2, respectively. All simulation data are available to download (Bower, 2025).
3.2.2 Volatile storage in the mantle
At fixed O2( IW = -3) and C/H (0.32 by mass; 100 solar), for low H mass fractions, CO(g) is the dominant species, but is exceeded by H2(g) across all hydrogen mass fractions 0.2% in the ideal case, whereas CH4 exceeds H2 above 0.3% H mass fraction in the real case. The predominance of CH4(g) results from the interplay between total pressure and the fugacity coefficients of the major atmospheric species, as described by the homogeneous equilibrium:
where, at equilibrium, and at constant O2:
where is the total mass of the atmosphere, is the acceleration due to gravity, and is the surface radius. Because pressure is proportional to 1/, even small changes in surface radius have tangible effects on total pressure, for a constant atmospheric mass budget. Conversely, the linear dependence of on (and hence on the planetary mass, ) results in a more modest mutual dependence. , and increasing pressure for a constant total mass budget results in higher dissolution (since , where is any given gas species), giving rise to the monotonic decrease in H2 (and hence total ) with increasing or with decreasing in the ideal case. Conversely, increasing leads to increases in (analogous to the mechanism discussed with respect to Equation 6, above), promoting additional H dissolution as molecular H2. The net result is sub-Neptunes are expected to be CH4-rich for cases in which mean densities are roughly 6500 kg/m3 at IW=-3, H=3% , and bulk C/H=100 solar. Figures G1 and G2 illustrate the effects of varying oxygen fugacity from (at constant C/H = 0.32 and H mass fraction = 1% ) and total C/H mass ratio (at constant O2 = IW=-3 and H mass fraction = 1% ), respectively, on volatile speciation in the atmosphere and dissolution in the interior assuming real and ideal gas behaviour, for K2-18b-like cases.
4 Discussion
4.1 Earth-sized planets
For Earth-sized planets, our results show that the dominant atmosphere at temperate conditions that emerges after planet formation is CO2-rich with a total surface pressure around 100 bar. Such an atmosphere forming upon cooling on the early Earth from a magma ocean state was first hypothesised by Holland (1984), computed by Zahnle et al. (2007) and demonstrated by Sossi et al. (2020) for a magma ocean initially equilibrated at its estimated O2 close to the IW buffer. CO2-rich atmospheres are compatible with that of Venus at the present day, both in composition and total pressure. The composition is also similar to that of modern Mars, hinting that this mode of atmosphere formation may have been commonplace on rocky planets in our solar system. That the Earth does not currently possess a 97:3 CO2–N2 atmosphere points to the operation of geochemical cycles that drew down carbon from the atmosphere through weathering by the Urey reaction to form carbonates (Sleep & Zahnle, 2001; Zahnle et al., 2007) that are subsequently recycled into the interior, thus evolving its atmosphere away from CO2-rich over geological timescales. This is because carbon is abundant, CO2 is strongly favoured at 280 K, and water condensation buffers the emergence of hydrogen-dominated atmospheres even when total C/H is low (as with atmospheres derived from a partially molten starting state, Appendix B).